# Diagnosing local minima and accelerating convergence of variational quantum eigensolvers with quantum subspace techniques

### v1

@@ Abstract
Recent research has shown that wavefunction evolution in real- and imaginary-time can generate quantum subspaces with significant utility for obtaining accurate ground state energies. Inspired by these methods, we propose combining quantum subspace techniques with the variational quantum eigensolver (VQE). In our approach, the parameterized quantum circuit is divided into a series of smaller subcircuits. The sequential application of these subcircuits to an initial state generates a set of wavefunctions that we use as a quantum subspace to obtain high-accuracy groundstate energies. We call this technique the circuit subspace variational quantum eigensolver (CSVQE) algorithm. By benchmarking CSVQE on a range of quantum chemistry problems, we show that it can achieve significant error reduction compared to conventional VQE, particularly for poorly optimized circuits, greatly improving convergence rates. Furthermore, we demonstrate that when applied to circuits trapped at a local minima, CSVQE can produce energies close to the global minimum of the energy landscape, making it a potentially powerful tool for diagnosing local minima.
@@ I Introduction
A main goal of quantum algorithms for Hamiltonian simulation is generating high fidelity groundstates for problems that are difficult to simulate classically. Two important classes of promising algorithms for achieving quantum advantage in Hamiltonian simulation are variational quantum algorithms [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 6, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] and quantum subspace methods [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. Variational quantum algorithms are widely studied, especially the variational quantum eigensolver (VQE), but quantum subspace methods are only recently beginning to grow in popularity. Quantum subspace methods have the potential to provide significant convergence improvements, but challenges remain in the efficient generation of these subspaces on quantum hardware.
In this work, we consider the use of VQE circuits as a mean for generating subspaces with accelerated convergence for groundstate energies of Hamiltonians. We draw inspiration for our approach from recent work studying subspaces generated using real- and imaginary-time evolution [76, 67, 77, 78]. We treat the VQE circuit itself as a time evolution operator in spirit, in which the application of each gate in the circuit acts like a discrete time evolution of the initial state. The initial state fed into the VQE circuit traces out a path through the Hilbert space as it is “time-evolved” by the VQE circuit, and it is from this path that we sample states to generate a subspace. For an optimized VQE circuit, one might informally consider the series of states generated by the circuit evolution to have similar characteristics to imaginary time evolution, as both rapidly evolve an initial state to the groundstate. By projecting the Hamiltonian for which the VQE circuit is optimized into this circuit-generated subspace, we form a generalized eigenvalue problem for the Hamiltonian. We call this method the circuit subspace variational quantum eigensolver (CSVQE). In principle, the subspace contains more information than the final state output by the VQE circuit, and the goal of this work is to quantify how much advantage can be extracted from this additional information.
II Circuit evolution subspace VQE
@@
We treat the parameters of these gates as constants throughout the CSVQE algorithm, assuming them to be determined by some classical optimization procedure. First we must generate the set of states explored by the circuit as it evolves the initial wavefunction via the sequential application of the quantum gates. Each of these states is obtained by the ordered application of the first $M\leq N$ gates in the circuit to the initial wavefunction,
@@
can be solved via classical algorithms for approximate eigenvalues of the original Hamiltonian. We remark that the overlap matrix, $S$, can be rank deficient, as is commonly encountered in other work on quantum subspace methods [67, 79]. This occurs most often when the mid-circuit states from nearby layers of the circuit have large overlaps with each other. In this work, we address this problem by projecting out singular values of $S$ that are less than some threshold, which we take to be $1\times 10^{-10}$. However, this threshold has to be adjusted in general when accumulating data from noisy hardware, as discussed in previous work [67]. The final wavefunction produced by VQE is limited by the ansatz and initial state chosen and, in most cases, is not capable of producing the exact groundstate. Solving the generalized eigenvalue problem utilizes these subspace states to generate a new groundstate wavefunction that has equal or larger overlap with the true groundstate, and also a lower energy than any individual wavefunction in the subspace.
@@
Figure 2: The error in the ground state energy produced by VQE and CSVQE at each step in the classical optimization process for LiH, $\mathrm{CH}_{4}$, $\mathrm{H}_{2}\mathrm{O}$, $\mathrm{NH}_{3}$, $\mathrm{N}_{2}$, and $\mathrm{C}_{2}$ using the UCC ansatz. The energies produced by the CSVQE algorithm using four, twelve, and all $(M=N+1)$ mid-circuit states are plotted in green squares, blue pentagons, and gray triangles, respectively. The reported errors represent the optimal values obtained by applying the CSVQE algorithm with 1000 different random combinations of mid-circuit states.
@@
As shown by the results in Fig. 1, the “even” and “front-loaded” strategies are comparably effective. Both reduce the error but do not quite achieve the accuracy of the best result obtained from 20,000 distinct randomly sampled state combinations. The “back-loaded” strategy obtains only minimal advantage over the standard VQE result. The intuition behind these results is as follows. We have ordered the gates in the circuit based on heuristic classical approximations, such that the gates that make the largest changes to the wavefunction are applied first in our construction of the circuit, which we discuss in detail in the following section. States obtained early in the circuit are therefore both the most distinct from the final state produced by the circuit, and also more distinct from their neighboring states than those later in the circuit. As such, they likely provide more utility when solving for the groundstate than the states at the end of the circuit, which have very large overlaps and contain nearly identical information. This is especially true in the case that circuit optimization obtains an excited state with only a small projection onto the groundstate, for which states near the beginning of the circuit likely contain much more useful overlaps with the groundstate.
@@ III Quantum chemistry benchmarking of CSVQE
$\begin{gathered}\hat{U}^{ab\cdots}_{ij\cdots}=\exp(\theta_{ij\cdots}^{ab\cdots}(\hat{a}_{ij\cdots}^{ab\cdots}\cdots-\hat{a}_{ab\cdots}^{ij\cdots}))\\ \hat{a}_{ij\cdots}^{ab\cdots}=\hat{a}^{\dagger}_{a}\hat{a}^{\dagger}_{b}\dots\hat{a}_{j}\hat{a}_{i}.\end{gathered}$ (12)
@@
    The circuits for each molecular are restricted to include at most fifty doubles operators, choosing the ones with the largest initial parameter values.
@@
    The circuit parameters are optimized using the BFGS implementation from SciPy, and we store the circuit obtained at each step of the optimization process [89].
@@
The reduction in error accomplished by the CSVQE algorithm ranges widely between both different molecules and different optimization steps of a given circuit. For all molecules studied here, CSVQE accomplishes significant error reduction by up to an order of magnitude early in the optimization process, often achieving parity with the error produced by the fully optimized circuit. However, the error reduction produced by the CSVQE algorithm diminishes as the optimization proceeds, indicating that optimized circuits produce wavefunctions with energies that cannot be improved via superposition with mid-circuit states (at least within our numerical tests).
@@ V Conclusions
In this work we devised a modification of the VQE algorithm, dubbed circuit-evolution VQE, that utilizes quantum spaces generated by sampling mid-circuit states from VQE circuits. We applied the CSVQE algorithm to UCC circuits for a range of molecules, studying the reduction in error that can be accomplished with the technique at each step in the optimization process. We found that the CSVQE algorithm accomplished significant reduction in error when applied to partially optimized circuits, but the advantage gained is minimal for fully optimized circuits. However, due to shot noise, hardware noise and difficult optimization landscapes, optimization on quantum hardware is likely to be less than ideal and the tools laid out here provide another route to improving accuracy in such cases.
@@
The utility of this algorithm is likely highly ansatz-dependent, and further investigation of its utility beyond the UCC ansatz and quantum chemistry problems is justified. One important shortcoming of the UCC ansatz utilized in this work is that a large portion of the UCC parameters are quite small, leading to a circuit evolution subspace containing many nearly-identical wavefunctions. The CSVQE algorithm will likely perform significantly better on more efficient ansatze in which each gate meaningfully changes the wavefunction. This would also significantly lessen the challenge of designing efficient state selection strategies for generating useful subspaces.
Acknowledgements.–We are grateful for support from NASA Ames Research Center. We acknowledge funding from the NASA ARMD Transformational Tools and Technology (TTT) Project. W.M. and N.T. acknowledge funding from the NASA ARMD Transformational Tools and Technology (TTT) Project. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359 (N.T.). M.R.H. participated in the Feynman Quantum Academy internship program. Y.S., K.K., and R.V.B. acknowledge funding from the U.S. Department of Energy (DOE) under Contract No. DE-AC0205CH11231, through the Office of Science, Office of Advanced Scientific Computing Research (ASCR) Exploratory Research for Extreme-Scale Science. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0024469.

### v2

# A circuit-generated quantum subspace algorithm for the variational quantum eigensolver

@@ Abstract
Recent research has shown that wavefunction evolution in real- and imaginary-time can generate quantum subspaces with significant utility for obtaining accurate ground state energies. Inspired by these methods, we propose combining quantum subspace techniques with the variational quantum eigensolver (VQE). In our approach, the parameterized quantum circuit is divided into a series of smaller subcircuits. The sequential application of these subcircuits to an initial state generates a set of wavefunctions that we use as a quantum subspace to obtain high-accuracy groundstate energies. We call this technique the circuit subspace variational quantum eigensolver (CSVQE) algorithm. By benchmarking CSVQE on a range of quantum chemistry problems, we show that it can achieve significant error reduction in the best case compared to conventional VQE, particularly for poorly optimized circuits, greatly improving convergence rates. Furthermore, we demonstrate that when applied to circuits trapped at a local minima, CSVQE can produce energies close to the global minimum of the energy landscape, making it a potentially powerful tool for diagnosing local minima.
@@ I Introduction
A main goal of quantum algorithms for Hamiltonian simulation is generating high fidelity groundstates for problems that are difficult to simulate classically. Two important classes of promising algorithms for achieving quantum advantage in Hamiltonian simulation are variational quantum algorithms [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 6, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] and quantum subspace methods [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. Variational quantum algorithms are widely studied, especially the variational quantum eigensolver (VQE), but quantum subspace methods are only recently beginning to grow in popularity. Quantum subspace methods have the potential to provide significant convergence improvements, but challenges remain in the efficient generation of these subspaces on quantum hardware.
In this work, we consider the use of VQE circuits as a mean for generating subspaces with accelerated convergence for groundstate energies of Hamiltonians. We draw inspiration for our approach from recent work studying subspaces generated using real- and imaginary-time evolution [79, 67, 80, 81]. We forgo the use of a Hamiltonian to guide the time evolution, and instead treat the VQE circuit itself as an evolution operator, in which the application of each gate in the circuit acts like a discrete evolution of the state, although not a proper real- or imaginary-time evolution. The initial state fed into the VQE circuit traces out a path through the Hilbert space as it is “evolved” by the VQE circuit, and it is from this path that we sample states to generate a subspace. By projecting the Hamiltonian for which the VQE circuit is optimized into this circuit-generated subspace, we form a generalized eigenvalue problem for the Hamiltonian. We call this method the circuit subspace variational quantum eigensolver (CSVQE). In principle, the subspace contains more information than the final state output by the VQE circuit, and the goal of this work is to quantify how much advantage can be extracted from this additional information.
II Circuit subspace VQE
@@
Here we define the gates $U_{i}$ as generic parameterized unitary operators, and in our numerics below they take the form of exponentiated antisymmetric combinations of fermionic excitation and deexcitation operators. For many ansatz it may be more natural to define $U_{i}$ as elementary quantum gates, and this requires no modification of the algorithm. In this work we take the parameters of the gates to be constants determined by some prior classical optimization procedure. The first step of the CSVQE algorithm is generating the set of states explored by the circuit as it evolves the initial wavefunction via the sequential application of the quantum gates. Each of these states is obtained by the ordered application of the first $M\leq N$ gates in the circuit to the initial wavefunction,
@@
can be solved via classical algorithms for approximate eigenvalues of the original Hamiltonian. We treat the circuit parameters as fixed in the application of the CSVQE algorithm, thus the CSVQE algorithm can be considered a post-processing step to be performed after a standard VQE calculation. The construction and solution of the above generalized eigenvalue problem can, in principle, be integrated into the optimization process of the VQE circuit, but we leave the question of how best to do that to future work.
We remark that the overlap matrix, $S$, can be rank deficient, as is commonly encountered in other work on quantum subspace methods [67, 82]. This occurs most often when the mid-circuit states from nearby layers of the circuit have large overlaps with each other. In this work, we address this problem by projecting out singular values of $S$ that are less than some threshold, which we take to be $1\times 10^{-10}$. However, this threshold has to be adjusted in general when accumulating data from noisy hardware, as discussed in previous work [67]. The final wavefunction produced by VQE is limited by the ansatz and initial state chosen and, in most cases, is not capable of producing the exact groundstate. Solving the generalized eigenvalue problem utilizes these subspace states to generate a new groundstate wavefunction that has equal or larger overlap with the true groundstate, and also a lower energy than any individual wavefunction in the subspace.
@@
Figure 2: The error in the ground state energy produced by VQE and CSVQE at each step in the classical optimization process for (a) LiH, (b) $\mathrm{CH}_{4}$, (c) $\mathrm{H}_{2}\mathrm{O}$, (d) $\mathrm{NH}_{3}$, (e) $\mathrm{N}_{2}$, and (f) $\mathrm{C}_{2}$ using the UCC ansatz. The energies produced by the CSVQE algorithm using four, twelve, and all $(M=N+1)$ mid-circuit states are plotted in green squares, blue pentagons, and gray triangles, respectively. The reported errors represent the optimal values obtained by applying the CSVQE algorithm with 1000 different random combinations of mid-circuit states.
@@
As shown by the results in Fig. 1, the “even” and “front-loaded” strategies are comparably effective. Both reduce the error but do not quite achieve the accuracy of the best result obtained from 20,000 distinct randomly sampled state combinations. The “back-loaded” strategy obtains only minimal advantage over the standard VQE result. Repeating these calculations for all other systems we study produces similar qualitative behavior for each state-selection strategy. The intuition behind these results is as follows. We have ordered the gates in the circuit based on heuristic classical approximations, such that the gates that make the largest changes to the wavefunction are applied first in our construction of the circuit, which we discuss in detail in the following section. States obtained early in the circuit are therefore both the most distinct from the final state produced by the circuit, and also more distinct from their neighboring states than those later in the circuit. As such, they likely provide more utility when solving for the groundstate than the states at the end of the circuit, which have very large overlaps and contain nearly identical information.
@@ III Quantum chemistry benchmarking of CSVQE
$\begin{gathered}\hat{U}^{ab\cdots}_{ij\cdots}=\exp(\theta_{ij\cdots}^{ab\cdots}(\hat{a}_{ij\cdots}^{ab\cdots}-\hat{a}_{ab\cdots}^{ij\cdots}))\\ \hat{a}_{ij\cdots}^{ab\cdots}=\hat{a}^{\dagger}_{a}\hat{a}^{\dagger}_{b}\dots\hat{a}^{\phantom{\dagger}}_{j}\hat{a}^{\phantom{\dagger}}_{i}.\end{gathered}$ (12)
@@
    The circuits for each molecule are restricted to include at most fifty doubles operators, choosing the ones with the largest initial parameter values.
@@
    The circuit parameters are optimized using the Broyden–Fletcher–Goldfarb–Shanno implementation from SciPy, and we store the circuit obtained at each step of the optimization process [92].
@@
The reduction in error accomplished by the CSVQE algorithm ranges widely between both different molecules and different optimization steps of a given circuit. For all molecules studied here, CSVQE accomplishes significant error reduction by up to an order of magnitude early in the optimization process, often achieving parity with the error produced by the fully optimized circuit. However, we note that this is an absolute best case scenario and practical experiment likely cannot achieve the number of random samples required to obtain these results. Furthermore, the error reduction produced by the CSVQE algorithm diminishes as the optimization proceeds, indicating that optimized circuits produce wavefunctions with energies that cannot be improved via superposition with mid-circuit states (at least within our numerical tests).
@@ V Conclusions
In this work we devised a modification of the VQE algorithm, dubbed circuit-subspace VQE, that utilizes quantum spaces generated by sampling mid-circuit states from VQE circuits. We applied the CSVQE algorithm to UCC circuits for a range of molecules, studying the reduction in error that can be accomplished with the technique at each step in the optimization process. We found that the CSVQE algorithm accomplished significant reduction in error in the best case when applied to partially optimized circuits, but the advantage gained is minimal for fully optimized circuits. However, due to shot noise, hardware noise and difficult optimization landscapes, optimization on quantum hardware is likely to be less than ideal and the tools laid out here provide another route to improving accuracy in such cases.
@@
The utility of this algorithm is likely highly ansatz-dependent, and further investigation of its utility beyond the UCC ansatz and quantum chemistry problems is justified. One important shortcoming of the UCC ansatz utilized in this work is that a large portion of the UCC parameters are quite small, leading to a circuit-generated subspace containing many nearly-identical wavefunctions. The CSVQE algorithm will likely perform significantly better on more efficient ansatze in which each gate meaningfully changes the wavefunction. This would also significantly lessen the challenge of designing efficient state selection strategies for generating useful subspaces. Another approach to easing the difficulty of choosing a useful subspace is to employ some measure of the overlap between the intermediate states as a selection criteria. The decoupling of the VQE optimization and the subspace calculations is a further area in which the algorithm could be improved, but we leave the question of how best to integrate circuit optimization with subspace calculations to future work.
We note that our approach has standard scaling with similar Krylov method techniques that require matrix elements to be measured on quantum hardware. However, the utility of the algorithm heavily depends on the quality of the optimization of the circuit, as demonstrated in the different case studied in this work. This is certainly of interest, as optimization on quantum hardware is resource intensive and will likely remain resource intensive even as quantum hardware improves. The above considerations imply that this algorithm may benefit significantly from integration with existing techniques [10, 15, 95].
Acknowledgements.–We are grateful for support from NASA Ames Research Center. We acknowledge funding from the NASA ARMD Transformational Tools and Technology (TTT) Project. W.M. and N.T. acknowledge funding from the NASA ARMD Transformational Tools and Technology (TTT) Project. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359 (N.M.T.). M.R.H. participated in the Feynman Quantum Academy internship program. Y.S., K.K., and R.V.B. acknowledge funding from the U.S. Department of Energy (DOE) under Contract No. DE-AC0205CH11231, through the Office of Science, Office of Advanced Scientific Computing Research (ASCR) Exploratory Research for Extreme-Scale Science. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0024469.